OpenStudy (anonymous):

which of the following is equal to the expression below ? (7^-5 ) 3 a. (-1/7 ) -15 b. (1/7 ) ^-15 c. (1/7) ^15 d.(-1/7 ) ^15 ? help plz thank u i appreciate it ????

OpenStudy (anonymous):

You didn't give me a medal for my answer last time :(

OpenStudy (anonymous):

help ??

OpenStudy (anonymous):

Lol okay. let me work this out

jimthompson5910 (jim_thompson5910):

(7^(-5))^3 = 7^(-5*3) = 7^(-15) = (7/1)^(-15) = (1/7)^(15) So the answer is (1/7)^(15)

OpenStudy (anonymous):

(7^(-5))^3 = 7^(-5*3) = 7^(-15) = (7/1)^(-15) = (1/7)^(15) So its (1/7)^(15)